5. The cross-price elasticity of demand is the percentage change in quantity divided by the
percentage change in price of another good.
6. The principle of diminishing marginal utility states that people enjoy consuming more of a
good.
7. If marginal utility is declining but still positive, total utility is decreasing.
8. ]n the long run all inputs are variable; in the short run all inputs are fixed.
9. Fixed costs decrease as output is increased.
10. The law of diminishing marginal productivity states that provided one input is constant,
increases in other inputs will eventually result in smaller increments to total output.
11. If total cost is 150, total fixed cost is 50, and output is 20, then average variable cost is 5.
12. An economically efficient method of production is a method that produces a given level of
output at the lowest possible cost.
13. In perfect competition, price is greater than marginal revenue.
14. An increase in the number of firms causes market supply to shift to the left.
15. A monopolistically competitive industry is characterized by few sellers In a highly
competitive market.
16. The demand for labor is based on the worker's trade-off between income and leisure.
Answer #1

5. True

The cross-price elasticity of demand is the percentage change in quantity divided by the percentage change in price of another good.

6. False

According to principle of diminishing marginal utiltiy, the marginal utility keeps on decreasing as consumption of one good keeps on increasing. Thus the principle of diminishing marginal utility states that people do not enjoy consuming more of a good.

7. False

Total utility is summation of marginal utility. If marginal utility is declining but still positive, total utility is increasing. Total utility starts decreasing when marginal utility is declining and negative.

8. False

In the long run all inputs are variable; in the short run some inputs are fixed and some are variable.

9. False

Fixed costs remain constant as output is increased. Average Fixed Cost decreases as output is increased.

10. True

The law of diminishing marginal productivity states that provided one input is constant, increases in other inputs will eventually result in smaller increments to total output.

11. True

Total cost= total fixed cost + total variable cost. Thus total variable cost= 150-50 = 100. Average variable cost = total variable cost / output = 100/20 = 5. Thus if total cost is 150, total fixed cost is 50, and output is 20, then average variable cost is 5.

12. True

An economically efficient method of production is a method that produces a given level of output at the lowest possible cost.

13. False

In perfect competition, price is equal to marginal revenue.

14. False

An increase in the number of firms causes market supply to shift to the right.

15. True

A monopolistically competitive industry is characterized by few sellers In a highly competitive market.

16. False

The supply of labor is based on the worker's trade-off between income and leisure.